determine the 8th term of the following geometric sequence?
3,9,27
write your answer as a single integer value, not an expression.



Answer by venugopalramana(3286)

determine the 8th term of the following geometric sequence?
3,9,27
write your answer as a single integer value, not an expression.
this is a gp with
a=3
r=3
n=8
tn=a*r^(n-1)=3*3^7= 2187